What is the range of function 3|sinx| - 4|cosx|? Please ellaborate the explanation..

Grade:

2 Answers

Nitish Jangid
11 years ago

we can get lowest value of sin = 0 , we can get higest value of sin = 1

we can get lowest value of cos = 0 , we can get higest value of cos = 1

when we put sins higest value then we gets coss lowest value..

so now we are putting the higest values of sin and lowest value of cos in function then ...

3|1| - 4|0| = 3-0 = 3

so now we are putting the higest values of cos and lowest value of sin in function then ...

3|0| - 4|1| = 0-4 = -4                   

so the range of function 3|sinx| - 4|cosx| is {-4 to 3}

Aniket Patra
11 years ago

Dear Raktim,

Nitish has solved it correctly.. Remmember the range of a function is its maximum and minimum possible values that can be attained within its domain of defination..

Wishes!